OpenStudy (cutebunny12):

Yuri solved the equation below incorrectly. Explain the step where he went wrong. fifteen times b minus one equals thirty Step 1: fifteen times b minus one minus one equals thirty minus one Step 2: fifteen times b equals twenty-nine Step 3: fifteen times b divided by fifteen equals twenty-nine divided by fifteen Step 4: b equals twenty-nine fifteenths

OpenStudy (anonymous):

15b - 1 = 30 15b - 1 + 1 = 30 15b = 31 15b / 15b = 31 / 15 \[b \approx 2.06\] Yuri did step 2 wrong, she subtracted one instead of adding it to isolate b.

OpenStudy (lynfran):

no he made a mistake at the very first step...

OpenStudy (cutebunny12):

thank u